OpenStudy (anonymous):

The fact that sin(150°) = .5 and sin(-150°) = -.5 implies that the sine function is which type of function?

OpenStudy (anonymous):

A) odd B) even C) neither D) cannot be determined

zepdrix (zepdrix):

even functions have this property: f(-x) = f(x) odd functions have this property: f(-x) = -f(x)

zepdrix (zepdrix):

If we take this first equation they gave us, multiply it by -1, then we get these two equations,\[\large\rm -\sin(150)=-\frac12,\qquad\qquad\qquad \sin(-150)=-\frac12\] They both equal -1/2, which tells us that\[\large\rm -\sin(150)=\sin(-150)\]

zepdrix (zepdrix):

So which property does it look like? Odd or even?

OpenStudy (anonymous):

even

zepdrix (zepdrix):

Ok let's make sure. If the sine function is our f, Then we have -f(150) = f(-150) We could call this 150 our x to simplify it further, -f(x) = f(-x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It's not even

zepdrix (zepdrix):

`odd functions have this property: f(-x) = -f(x)` that's what we ended up with, right?

OpenStudy (anonymous):

yes